Find the power or energy of the following signals. xt=0                                 &nb

  1. Find the power or energy of the following signals.

  1. xt=0                                          for t<0(e-t-e-10t)                 for 0 ≤ t<2 (1-e-5)e-t                       for t≥2       (16 points)

  1. xt=6+5cos30t+3cos60t                              (8 points)
Answer #1

solno Criven, a alt= & Lo [pager) let ost 2 e lot) (1-eslet of ar Energy of alta sincetat E[nt] = I (et e 10t)? 2 ettt S6-e5epage-2 (6) r(t) = 6 + 5 cassot + 3 cos fot rus (a(H)] = √ (612+ ( 1 ) + ( 2 ) sus (DC) - DC sus (Ani Cosent) = Any 2 - 2 + 2